OpenStudy (anonymous):

Which of the following are measurements of the sides of a right triangle? 101, 99, 20 28, 26, 12 17, 14, 6 Which of the following are measurements of the sides of a right triangle? 101, 99, 20 28, 26, 12 17, 14, 6 none of the above none of the above

OpenStudy (anonymous):

sorry I wrote it twice

OpenStudy (cwrw238):

check to see if they fit the pythagoras theorem eg is 101^2 = 100^2 + 20^2 ?

OpenStudy (cwrw238):

sorry is 101^2 = 99^2 + 20^2 ?

OpenStudy (cwrw238):

if so it is right angled

OpenStudy (triciaal):

for a right triangle the hypotenuse is the longest side and opposite the right angle. then plug the numbers correctly to follow what @cwrw238 wrote

OpenStudy (anonymous):

thanks you guys

OpenStudy (triciaal):

you are welcome

OpenStudy (anonymous):

so why do you guys do this ? help us not sure people all day everyday

OpenStudy (cwrw238):

well believe it or not i like doing it - i enjoy math

OpenStudy (cwrw238):

though i'm not doing it all day!!

OpenStudy (anonymous):

I understand I enjoy certain parts of math I love solving for x or midpoint ect lol hey you never now

OpenStudy (anonymous):

or anything I can just plug in to a formula or calculator

OpenStudy (triciaal):

fun, easy, relaxing, helping students to succeed, educational, safe, etc

OpenStudy (anonymous):

good for you =) but I do not know how math could be dangerous haha im just joking sorry

OpenStudy (triciaal):

Math was a part of me. I found this site and now it is an addiction. To help others with just the little extra he or she may need is very rewarding.

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

its very rewarding

OpenStudy (anonymous):

:) im more of a history know it but i understand I bet they all thank you

OpenStudy (anonymous):

*knew it all

OpenStudy (triciaal):

all the best in your studies. history never changes, more stories to tell, continuous

OpenStudy (anonymous):

thank you